# Unveiling Entity-Level Unlearning for Large Language Models: A Comprehensive Analysis

### v2

@@ Abstract
Large language model unlearning has garnered increasing attention due to its potential to address security and privacy concerns, leading to extensive research in the field. However, much of this research has concentrated on instance-level unlearning, specifically targeting the removal of predefined instances containing sensitive content. This focus has left a significant gap in the exploration of full entity-level unlearning, which is critical in real-world scenarios such as copyright protection. To this end, we propose a novel task of Entity-level unlearning, which aims to erase entity-related knowledge from the target model completely. To thoroughly investigate this task, we systematically evaluate trending unlearning algorithms, revealing that current methods struggle to achieve effective entity-level unlearning. Then, we further explore the factors that influence the performance of the unlearning algorithms, identifying that knowledge coverage and the size of the forget set play pivotal roles. Notably, our analysis also uncovers that entities introduced through fine-tuning are more vulnerable to unlearning than pre-trained entities. These findings collectively offer valuable insights for advancing entity-level unlearning for LLMs.
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a, b; Meta, 2024) pre-trained on extensive corpora have achieved significant success in knowledge-intensive tasks (Kamalloo et al., 2023; Seegmiller et al., 2024). However, undesirable data exists with training data, such as toxic texts (Lu et al., 2022), privacy content (Liu et al., 2024a) and copyrighted information (Karamolegkou et al., 2023). Such data has raised security and legal concerns, hindering the practical application of LLMs (Yao et al., 2024; Das et al., 2024). To tackle this, Machine Unlearning (Zhang et al., 2023; Lu et al., 2024; Bhardwaj et al., 2024) has gradually been applied to LLMs due to its effectiveness and cost-efficiency. These refined techniques, now known as LLM Unlearning (Yao et al., 2023; Liu et al., 2024b, c), have become a mainstream approach for removing undesirable knowledge from the model by applying post-hoc modifications to target models.
@@
Figure 1: The comparison of the entity-level unlearning and instance unlearning processes. The knowledge covered by the purple background represents the target set, and the knowledge covered by the green background represents the forget set.
@@
Towards this direction, existing work has conducted extensive research. However, most of these efforts focus on Instance-level Unlearning tasks, which address isolated sensitive content (Li et al., 2024; Zhang et al., 2024; Ji et al., 2024), while neglecting the deletion of entire entities, which is crucial in many real-world scenarios, such as removing ‘Harry Potter’ for copyright protection Eldan and Russinovich (2024). To address this gap, we formally define a novel task of Entity-level Unlearning. As illustrated in Figure 1, existing instance-level unlearning tasks focus on removing predefined facts by applying unlearning algorithms to a forget set that contains the specific information to be erased. In contrast, the entity-level unlearning task seeks to eliminate an entire entity, including all the entity-related knowledge within the model. However, due to the inaccessibility of the original training corpora, it is challenging to obtain such a target set, which contains all knowledge associated with the target entity. Consequently, entity-level unlearning requires an additional step to construct a forget set before applying unlearning algorithms, where there is frequently a discrepancy between the constructed forget set and the target set. To this end, we divide the entity-level unlearning task into a two-stage framework, comprising forget set construction and unlearning execution. This process seeks to delete the target set by erasing the construted forget set, which involves more complex procedures and presents greater challenges than instance-level unlearning.
In this work, we conduct a comprehensive analysis of entity-level unlearning for LLMs, structured around a sequence of focused steps: Firstly, we begin by formally defining and setting the task, which involves injecting pseudo-entity knowledge into the model to establish both the target model and the target set. Secondly, we systematically evaluate the current unlearning algorithms on this task. Specifically, within the framework, we introduce an accessible knowledge-probing method to identify entity-related knowledge as the forget sets, and then compare the performance of the unlearning algorithm across different forget sets. Our experiments reveal that existing unlearning algorithms can not generalize well to the entity-level unlearning task. Additionally, we observe that the design of forget set significantly influences the effectiveness of the unlearning algorithm. Thirdly, inspired by the above, we further investigate how various constructed forget sets impact algorithm performance. Our results indicate that unlearning with a forget set that has higher knowledge coverage relative to the target set leads to more effective unlearning. Although merely increasing the size of the forget set may enhance knowledge coverage, it would compromise the model’s generalization ability, making it an ineffective way to improve performance. Moreover, our experiments surprisingly reveal that entities introduced through fine-tuning are more vulnerable to unlearning compared to pre-trained entities, highlighting the need for more robust methods for injecting target knowledge to simulate pre-trained knowledge better.
Through extensive experiments, our study provides valuable insights for future research. We encourage future work to explore more precise knowledge-probing techniques and develop targeted unlearning algorithms to improve the preformance of entity deletion.
@@ 2.1 Task Definition and Setting
The objective of the entity-level unlearning task is to remove an entire entity from the target model completely. However, the inaccessible training data makes it challenging to obtain comprehensive information about the entity within the model. Therefore, we define the task as removing the entire entity from the target model by deleting only a subset of entity-associated knowledge.
The entity-level unlearning task can be formalized as follows: Given a target entity O, the target model parametrized by $\theta_{t}$ is required to forget a target set $S_{T}$, which contains all knowledge related to the entity O in the model, by applying unlearning methods $\textsc{H}(\cdot)$ on a forget set $S_{F}$, which contains part of the knowledge of the entity. All of the aforementioned sets consist of pieces of knowledge, $S=\{(x_{i},y_{i})\}_{i=1}^{N}$. The unlearning process can be expressed as follows:
@@
A significant challenge in evaluating the effectiveness of entity-level unlearning posed by the inability to access all their training data of LLMs is obtaining the target set $S_{T}$ used to evaluate the unlearned models. To address this, we simulate entity-level unlearning scenarios following the TOFU (Maini et al., 2024), which fine-tunes the LLMs using a fictitious author dataset. The dataset ensures that the LLM has no prior exposure to these authors during previous training phases. Thus, the fine-tuning dataset encompasses all the knowledge about the entities, making it suitable to serve as the target set $S_{T}$.
@@ 2.2.1 Forget Set Construction
Specifically, we first prompt the target models to self-generate entity-related questions according to their internal knowledge, inspired by Weller et al. (2023). Note that only those non-repetitive questions containing the entities’ names will be kept to further ensure the focus remains on the corresponding entities. Next, we acquire the answers from the target model using greedy decoding, which ensures that the answers possess a relatively high generation probability. Finally, we conduct a self-verification process, where the model repeatedly evaluates each QA pair, retaining only the pairs for which the model consistently agrees with its responses. These selected pairs serve as valid candidates for forget sets of unlearning techniques.
@@ 2.2.2 Unlearning Execution
After constructing the forget set, the next step is to apply the unlearning algorithms to it. In the absence of algorithms specifically designed for entity-level unlearning, we select five representative algorithms for evaluation. Among them is Gradient Ascent (Grad. Ascent), which reduces the likelihood of the answers to achieve unlearning. Other methods introduce additional constraints on a retain set, which contains knowledge that should be preserved to minimize the damage to the model. An ideal entity-level unlearning method should effectively remove the entire entity from the forget set while minimizing any negative impact on the remaining knowledge. This study primarily examines the performance of current unlearning algorithms on entity-level unlearning tasks, without yet delving into the broader algorithmic application framework (Huang et al., 2024).
@@ 3.1 Experimental Setup
  • Target Set: For the target entity, we select the original training dataset with 20 question-answer pairs in TOFU as the target set.
  • Forget Set: In the experiments, we define two types of forget sets. The first type involves selecting the target set as the forget set for the unlearning algorithm, simulating an ideal scenario. The second type is constructed by generating a probing set from the target model, following the methods outlined in Section 2.2.1. The probing set, used as the forget set, consists of 20 QA pairs related to specific entities.
  • Retain Set: Unlike TOFU, we construct a retain set using questions with greedy-decoding answers from TriviaQA Joshi et al. (2017). These questions could be correctly answered by both Llama2-7B-Chat-TOFU and Phi-1.5-TOFU.
@@
  • ROUGE (Lin, 2004) measures the overlap co-occurrence of n-grams between the original answer and model’s greed-decoding generation for the test QA pairs.
@@ 3.2 Experimental Results
A comparison of the performance across different types of forget sets, as presented in Table 1, reveals that the algorithms based on the target set consistently maintain similar model utility while achieving lower probability, reduced accuracy, and overall higher forget quality on Llama2-7B-Chat-TOFU. A similar trend is observed in the Phi-1.5-TOFU model, as shown in Table 4. These results suggest that models employing the target set engage in more thorough unlearning compared to those utilizing the probing set. This finding highlights that the current unlearning algorithms struggle to generalize effectively to entity-level unlearning tasks when relying on the probing set. Consequently, the construction of the forget set is crucial in determining the success of entity-level unlearning. In the subsequent analysis, we will explore in detail how the quality of the forget set influences unlearning effectiveness.
@@
    The NPO-GD method stands out as one of the strongest approaches in unlearning, particularly excelling in instance-level unlearning tasks under ideal conditions. However, it exhibits the poorest performance on the probing set, with a significant disparity in forget quality between the target set and the probing set: 0.5253 vs. 0.0875 for Llama2-7B-Chat-TOFU and 0.5104 vs. 0.0794 for Phi-1.5-TOFU, according to Table 1 and 4. This substantial disparity may be attributed to the current optimization focus of unlearning algorithms, which tends to prioritize the forget set itself, often in opposition to the goal of generalizing within the entity, a crucial aspect for effective entity-level unlearning. This highlights the necessity for developing a specialized algorithm tailored specifically for entity-level unlearning.
@@ 4 Analysis
Figure 2: The performance of five unlearning algorithms in forget quality and model utility metrics across different probing sets, achieved by replacing varying ratios of QA pairs from the target set.
@@
The main experimental results indicate that the unlearning algorithms exhibit limited effectiveness in removing entities with the probing set. In this section, we analyze the pivot factors that may influence the effect of the algorithms. Our subsequent analysis is conducted on the Llama2-7B-Chat-TOFU model, which is widely used.
4.1 Effect of Probing Set
@@
We hypothesize that the limited unlearning effect is due to the insufficient coverage of knowledge within the current probing set. To explore this, we introduce a new metric, Knowledge Coverage, designed to assess the knowledge overlap between the probing sets and the target set. This metric is computed using the BERTScore Zhang et al. (2019) of the closest QA pair match between the probing set and the target set. A higher knowledge coverage indicates that the probing set encompasses a broader range of entity-related knowledge. See Appendix C.1 for detailed definitions and formulas.
@@
To further investigate the effects of varying knowledge coverage, we construct probing sets with varying degrees of coverage by systematically replacing different ratios of QA pairs within the probing set with those from the target set while keeping the total set size fixed at 20. As demonstrated in Table 7, the knowledge coverage of the constructed probing set increases progressively as the ratio of replaced pairs grows. Then, we applied the five algorithms to perform entity-level unlearning on the probing set under five different replacement ratios. As illustrated in Figure 2, enhancing knowledge coverage of the probing set consistently enhances the forget quality of most methods. Furthermore, each algorithm consistently preserved similar model utility across various probing sets. These results suggest that enhancing the knowledge coverage of the probing set can improve the unlearning effectiveness.
@@
An intuitive approach to increase knowledge coverage is by increasing the size of the probing set. Therefore, we further explore the performance of the algorithms on probing sets across different sizes. Specifically, we expand the probing set size for each entity and evaluate the performance of the unlearning algorithms across five different sizes. As illustrated in Figure 3, the forget quality of the five unlearning algorithms gradually improves as the size of the probing set increases. However, except for the Pref. Opt. method, the model utility of the unlearned models produced by the other algorithms demonstrates a noticeable decline. This decline in model utility, when compared to the trends in Figure 2, is likely due to the larger number of unlearning steps required as the probing set size grows. These findings suggest that, for most unlearning algorithms, expanding the probing set size leads to a trade-off between forget quality and model utility. Furthermore, according to section 3.2, although the Pref. Opt. method maintains a high generation probability for the forget knowledge, it fails to achieve true deletion, as the forgotten information can still be easily retrieved through fill-in-the-blank questions (Wang et al., 2024a).
@@
Figure 3: The performance of five unlearning algorithms in forget quality and model utility metrics under different scale probing sets.
@@
Furthermore, we also measure the knowledge coverage and record the probing rounds of probing sets across different sizes. As shown in Table 2, knowledge coverage truly increases with the size of the probing set, but the rate of increase gradually diminishes. Simultaneously, expanding the size of the probing set proves to be challenging. We employ the knowledge-probing method to iteratively enlarge the dataset and report the number of iterations required for different set sizes in Table 2. The results indicate that once the probing set reaches a certain size, further expansion of the probing set becomes increasingly resource-intensive. The analysis mentioned indicates that simply increasing the size of the probing sets does not effectively enhance the unlearning performance.
@@
Table 2: Knowledge Coverage across various probing set sizes and the average number of probing rounds required to obtain each set. Values in () indicate the increase relative to the previous set.
In summary, the quality of the probing set plays a crucial role in the unlearning effect. A probing set with greater knowledge coverage results in better unlearning performance. However, using the large probing sets would compromise model utility and incur higher costs. Therefore, we encourage future research to explore more precise knowledge probing methods and to construct higher-quality probing sets as forget sets.

### v3

@@ Abstract
Large language model unlearning has garnered increasing attention due to its potential to address security and privacy concerns, leading to extensive research in the field. However, existing studies have predominantly focused on instance-level unlearning, specifically targeting the removal of predefined instances containing sensitive content. This focus has left a gap in the exploration of removing an entire entity, which is critical in real-world scenarios such as copyright protection. To close this gap, we propose a novel task named Entity-level unlearning, which aims to erase entity-related knowledge from the target model completely. To investigate this task, we systematically evaluate popular unlearning algorithms, revealing that current methods struggle to achieve effective entity-level unlearning. Then, we further explore the factors that influence the performance of unlearning algorithms, identifying that the knowledge coverage of the forget set and its size play pivotal roles. Notably, our analysis also uncovers that entities introduced through fine-tuning are more vulnerable than pre-trained entities during unlearning. We hope these findings can inspire future improvements in entity-level unlearning for LLMs.
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a, b; Meta, 2024) pre-trained on extensive corpora have achieved significant success in various downstream tasks (Kamalloo et al., 2023; Seegmiller et al., 2024). However, training data often contains undesirable information, such as toxic texts (Lu et al., 2022), privacy content (Liu et al., 2024a) and copyrighted information (Karamolegkou et al., 2023). These issues raise security and legal concerns, hindering the practical application of LLMs (Yao et al., 2024; Das et al., 2024). To tackle this, Machine Unlearning (Zhang et al., 2023; Lu et al., 2024; Bhardwaj et al., 2024) has gradually been applied to LLMs due to its effectiveness and cost-efficiency. These refined techniques, now known as LLM Unlearning (Yao et al., 2023; Liu et al., 2024b, c), have become a mainstream approach for removing undesirable knowledge from the model by applying post-hoc modifications to target models.
@@
Figure 1: The comparison between the instance-level unlearning process and entity-level unlearning process. The knowledge covered by the purple background represents the target set, and the knowledge covered by the green background represents the forget set.
@@
Towards this direction, the current unlearning paradigm involves applying an algorithm on a forget set which contains the undesirable knowledge (Zhang et al., 2024; Maini et al., 2024). However, most existing research focuses on Instance-level Unlearning tasks, which address isolated sensitive content (Li et al., 2024; Ji et al., 2024), while overlooking the deletion of entire entities, which is crucial in many real-world scenarios, such as removing ‘Harry Potter’ for copyright protection Eldan and Russinovich (2024). To address this gap, we formally define a novel task named Entity-level Unlearning. As illustrated in Figure 1, this task represents a significant divergence from instance-level unlearning due to the different nature of the information being targeted for removal. Specifically, instance-level unlearning focuses on removing specific facts, which are predefined and can be directly erased. In contrast, entity-level unlearning addresses the removal of an entire entity, requiring the deletion of all knowledge associated with the entity within the model, which is referred to as the target set. However, in real-world scenarios, the knowledge about an entity in the model is an abstract set that cannot be predefined. Consequently, entity-level unlearning requires an additional step to construct a specific forget set containing the entity-related knowledge for deletion.
To this end, we propose a two-stage framework for the task, comprising forget set construction and unlearning execution. Specifically, in the first stage, we introduce a knowledge-probing method to obtain entity-related knowledge as the forget set. In the second stage, we perform removing the target entity by applying an unlearning algorithm to the constructed forget set. To establish a controlled experimental environment, we choose to inject pseudo-entity knowledge into the model through the post-training, ensuring that all the knowledge of the target entity (target set) is accessible. This allows us to assess the unlearning effect of pseudo-entity on the target set, which can achieve a precise evaluation. Based on the settings, we systematically analyze the performance of five popular unlearning algorithms on the entity-level unlearning task within the two-stage framework and further investigate the pivot factors influencing the unlearning effect. Through extensive experiments, we have drawn several insightful conclusions:
  • Existing unlearning algorithms, which only focus on the forget set itself, struggle to achieve effective entity-level unlearning under real scenarios, making it necessary to develop corresponding unlearning algorithms that can generalize deletion effects to the entire entity’s knowledge.
  • Increasing the knowledge overlap between the constructed forget set and the target set can enhance the entity deletion effect of algorithms. However, simply expanding the size of the set is an ineffective way to improve this overlap, which will destroy the general ability of the model.
  • The entities introduced through fine-tuning are more vulnerable than the pre-trained entities during unlearning, highlighting the need for more robust methods for injecting target knowledge to simulate pre-trained knowledge better.
In summary, removing an entire entity from LLMs is a realistic and challenging task. We encourage future research to build on our findings by exploring more precise knowledge-probing methods and developing targeted unlearning algorithms to enhance the effectiveness of entity deletion.
@@ 2.1 Task Definition and Setting
The objective of the entity-level unlearning task is to remove an entire entity from the target model. However, the knowledge associated with an entity is an abstract concept, which cannot be predefined in real-world scenarios. Therefore, it is necessary to construct a specific set that contains the model’s knowledge of the entity for deletion. Consequently, we define this task as removing the entire entity from the target model by deleting a specific set containing entity-associated knowledge.
The entity-level unlearning task can be formalized as follows: Given a target entity O, the target model parametrized by $\theta_{t}$ is required to forget a target set $S_{T}$, which contains all knowledge related to the entity O in the model, by applying unlearning methods $\textsc{H}(\cdot)$ on a forget set $S_{F}$, which contains the specific knowledge of the entity. The unlearning process can be expressed as follows:
@@
A significant challenge in evaluating the effectiveness of entity-level unlearning posed by the inability to access all their training data of LLMs is obtaining the target set $S_{T}$ used to evaluate the unlearned models. To address this, we simulate entity-level unlearning scenarios following the TOFU (Maini et al., 2024), which fine-tunes the LLMs using a fictitious author dataset. The dataset ensures that the LLMs have no prior exposure to these authors during previous training phases. Thus, the fine-tuning dataset encompasses all the knowledge about the entities, making it suitable to serve as the target set $S_{T}$.
@@ 2.2.1 Forget Set Construction
Specifically, we first prompt the target models to self-generate entity-related questions according to their internal knowledge, inspired by Weller et al. (2023). Note that only those non-repetitive questions containing the entities’ names will be kept to further ensure the focus remains on the corresponding entities. Next, we acquire the answers from the target model using greedy decoding, which ensures that the answers possess a relatively high generation probability. Finally, we conduct a self-verification process, where the model repeatedly evaluates each QA pair, retaining only the pairs for which the model consistently agrees with its responses. These selected pairs serve as valid candidates for forget sets of unlearning algorithms.
@@ 2.2.2 Unlearning Execution
After constructing the forget set, the next step is to apply the unlearning algorithms to it. In the absence of algorithms specifically designed for entity-level unlearning, we select five representative algorithms for evaluation. One of the most straightforward unlearning methods is Gradient Ascent, which reduces the likelihood of the answers to achieve unlearning. Other methods introduce additional constraints on a retain set, which contains knowledge that should be preserved to minimize the damage to the model. An ideal entity-level unlearning method should effectively remove the entire entity while minimizing any negative impact on the remaining knowledge. This study primarily examines the performance of current unlearning algorithms on the entity-level unlearning task, without yet delving into the broader algorithmic application framework (Huang et al., 2024).
@@ 3.1 Experimental Setup
  • Target Set: For the target entity, we select the oracle training dataset with 20 question-answer pairs in TOFU as the target set.
  • Forget Set: In the experiments, we define two types of forget sets. The first type involves selecting the target set as the forget set for the unlearning algorithms, simulating an ideal scenario. The second type is constructed by generating a constructed forget set (Con. forget set) from the target model, following the methods outlined in Section 2.2.1. The constructed forget set consists of 20 QA pairs related to the target entity.
  • Retain Set: The retain set is constructed from TriviaQA Joshi et al. (2017) and consists of QA pairs related to world knowledge. We ensure that each question is correctly answered by both two target models using greedy decoding.
@@
  • ROUGE (Lin, 2004) measures the overlap co-occurrence of n-grams between the original answer and the model’s greed-decoding generation for the test QA pairs.
@@ 3.2 Experimental Results
A comparison of the performance across different types of forget sets, as presented in Table 1, reveals that the algorithms based on the target set consistently maintain similar model utility while achieving lower probability, reduced accuracy, and overall higher forget quality on Llama2-7B-Chat-TOFU. A similar trend is observed in the Phi-1.5-TOFU model, as shown in Table 4. These results suggest that applying the unlearning algorithms to the target set results in more thorough forgetting than their application on the constructed forget set. This finding highlights that the current unlearning algorithms struggle to generalize effectively to entity-level unlearning tasks when relying on the constructed forget set. Consequently, the construction of the forget set is crucial in determining the success of entity-level unlearning. In the subsequent analysis, we will explore in detail how the quality of the forget set influences unlearning effectiveness.
@@
    The NPO-GD method stands out as one of the strongest approaches in unlearning, particularly excelling in instance-level unlearning tasks under ideal conditions. However, it exhibits the poorest performance on the constructed forget set, with a significant disparity in forget quality between the target set and the constructed forget set: 0.5253 vs. 0.0875 for Llama2-7B-Chat-TOFU and 0.5104 vs. 0.0794 for Phi-1.5-TOFU, according to Table 1 and 4. This substantial disparity may be attributed to the current optimization focus of unlearning algorithms, which tends to prioritize the forget set itself, often in opposition to the goal of generalizing within the entity, a crucial aspect for effective entity-level unlearning. This highlights the necessity for developing a targeted algorithm tailored specifically for entity-level unlearning.
@@ 4 Analysis
Figure 2: The performance of five unlearning algorithms in forget quality and model utility metrics across different constructed forget sets, achieved by replacing varying ratios of QA pairs from the target set.
@@
The main experimental results indicate that the unlearning algorithms exhibit limited effectiveness in removing entities with the constructed forget set. In this section, we analyze the pivot factors that may influence the effect of the algorithms. Our subsequent analysis is conducted on the Llama2-7B-Chat-TOFU model, which is widely used.
4.1 Effect of Forget Set
@@
We hypothesize that the limited unlearning effect is due to the insufficient coverage of knowledge within the current constructed forget set. To explore this, we introduce a new metric, Knowledge Coverage, designed to assess the knowledge overlap between the constructed forget sets and the target set. This metric is computed using the BERTScore Zhang et al. (2019) of the closest QA pair match between the constructed forget set and the target set. A higher knowledge coverage indicates that the constructed forget set encompasses a broader range of entity-related knowledge. See Appendix C.1 for detailed definitions and formulas.
@@
To further investigate the effects of varying knowledge coverage, we construct forget sets with varying degrees of coverage by systematically replacing different ratios of QA pairs within the constructed forget set with those from the target set while keeping the total set size fixed at 20. As demonstrated in Table 7, the knowledge coverage of the constructed forget set increases progressively as the ratio of replaced pairs grows. Then, we applied the five algorithms to perform entity-level unlearning on the constructed forget set under five different replacement ratios. As illustrated in Figure 2, enhancing knowledge coverage of the constructed forget set consistently enhances the forget quality of most methods. Furthermore, each algorithm consistently preserved similar model utility across various constructed forget sets. These results suggest that enhancing the knowledge coverage of the constructed forget set can improve the unlearning effectiveness.
@@
An intuitive approach to increase knowledge coverage is by increasing the size of the constructed forget set. Therefore, we further explore the performance of the algorithms on constructed forget sets across different sizes. Specifically, we expand the constructed forget set size for each entity and evaluate the performance of the unlearning algorithms across five different sizes. As illustrated in Figure 3, the forget quality of the five unlearning algorithms gradually improves as the size of the constructed forget set increases. However, except for the Pref. Opt. method, the model utility of the unlearned models produced by the other algorithms demonstrates a noticeable decline. This decline in model utility, when compared to the trends in Figure 2, is likely due to the larger number of unlearning steps required as the constructed forget set size grows. These findings suggest that, for most unlearning algorithms, expanding the constructed forget set size leads to a trade-off between forget quality and model utility. Furthermore, according to section 3.2, although the Pref. Opt. method maintains a high generation probability for the forget knowledge, it fails to achieve true deletion, as the forgotten information can still be easily retrieved through fill-in-the-blank questions (Wang et al., 2024).
@@
Figure 3: The performance of five unlearning algorithms on the constructed forget sets across different scales in forget quality and model utility metrics
@@
Furthermore, we also measure the knowledge coverage and record the probing rounds of constructed forget sets across different sizes. As shown in Table 2, knowledge coverage truly increases with the size of the constructed forget set, but the rate of increase gradually diminishes. Simultaneously, expanding the size of the constructed forget set proves to be challenging. We employ the knowledge-probing method to iteratively enlarge the dataset and report the number of iterations required for different set sizes in Table 2. The results indicate that once the constructed forget set reaches a certain size, further expansion of the constructed forget set becomes increasingly resource-intensive. The analysis mentioned indicates that simply increasing the size of the constructed forget sets does not effectively enhance the unlearning performance.
@@
Table 2: Knowledge Coverage across various constructed forget set sizes and the average number of probing rounds required to obtain each set. Values in () indicate the increase relative to the previous set.
In summary, the quality of the constructed forget set plays a crucial role in the unlearning effect. A constructed forget set with greater knowledge coverage results in better unlearning performance. However, using the large constructed forget sets would compromise model utility and incur higher costs. Therefore, we encourage future research to explore more precise knowledge probing methods to construct higher-quality constructed forget sets.